OpenStudy (anonymous):

What is the direct linear variation equation for the relationship? y varies directly with x and y = 4 when x = 12. A. y = 3x B. 01_15_UT_07a.gif C. y = x – 8 D. y = x + 8

OpenStudy (anonymous):

B.y=1/3x

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@openstudier101x @bohotness @dan815

OpenStudy (anonymous):

The equation for direct linear variation is y = kx. Your x and y coordinates are given as 4 and 12, so to put this into direct linear variation format you need to find k first. To do so, you substitute the equation with your given values: 4 = k(12) 4/12 = k 1/3 = k So, your final equation is y=1/3x, and you are not using your given values anymore because the final answer for direct linear variation is intended so that you could use this equation for any given point with any x or y value when needed.

OpenStudy (anonymous):

Happy studying! :)